OpenStudy (anonymous):

The value of 12 is obtained by multiplying p% by a-12 b-12/p c-p/12 d-1200/p e-100p/12

OpenStudy (anonymous):

let us do it this way: k*p%=k*p/100=12---->k=12x100/p=1200/p got it?

OpenStudy (anonymous):

Yerp!